OpenStudy (anonymous):

Let f(x) = x + 9 and g(x) = x – 3. What is f(x) • g(x)?

Parth (parthkohli):

\(f(x) \cdot g(x) = (x + 9)(x - 3)\)

Parth (parthkohli):

Simplify that

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x + 9)(x - 3) =x(x-3) + 9(x - 3)}\)

OpenStudy (anonymous):

Just multiply (x + 9) by (x - 3). (x + 9)(x - 3) x^2 - 3x + 9x - 27 x^2 + 6x - 27 That's your answer. :)

OpenStudy (unklerhaukus):

unless you ment \[f\circ g(x)\neq f(x)\cdot g(x)\]